Chewy Inc is a retailer operating in the town. Chewy uses the perpetual inventory method. All ales return from the customer result in the goods being returned to inventory, the inventory is not damage. Assume that there are no credit transactions; all amounts are settled in cash. You are provided with the following information for Chewy Inc. for the month of January 2018. Description Quantity Unit cost or selling price Date $15 14 25 25 January 1 Beginning inventory January 5 Purchase January 8 Sale January 10 Sales return January 15 Purchase January 16 Purchase return January 20 Sale January 25 Purchase 60 110 90 10 35 10 90 10 25 20 Required: (a) For each the following cost flow assumptions, calculate (i) cost of goods sold, (ii) ending inventory and (iii) gross profit. (1) LIFO cot (2) FIFO (3) Moving-average
Answer #1

1)Under LIFO Method, goods purchased latest are sold first

Transaction

Quantity

Rate

Amount

Jan1 Beginning Inventory

60

15

900

Jan5 Purchases

110

14

1,540

Jan8 Sale

(90)

14

(1,260)

Jan10 Sales Return

10

14

140

Jan15 Purchase

35

18

630

Jan16 Purchase Return

(10)

18

(180)

Jan20 Sales

(25)

18

(450)

Jan20 Sales

(30)

14

(420)

Jan20 Sales

(35)

15

(525)

Jan25 Purchase

10

20

200

Ending Inventory

35

575

Cost of Goods Sold = $1,260-140+450+420+525 = $2,515

Gross Profit :

Sales 80*25 +90*25 = $4,250

Less: COGS = $2,515

Gross Profit = $1,735

2)FIFO

Inventory purchased first is sold first

Transaction

Quantity

Rate

Amount

Jan1 Beginning Inventory

60

15

900

Jan5 Purchases

110

14

1,540

Jan8 Sale

(60)

15

(900)

Jan8 Sale

(30)

14

(420)

Jan10 Sales Return

10

14

140

Jan15 Purchase

35

18

630

Jan16 Purchase Return

(10)

18

(180)

Jan20 Sales

(90)

14

(1,260)

Jan25 Purchase

10

20

200

Ending Inventory

35

650

Cost of Goods Sold = $900+420-140+1,260 = $2,440

Gross Profit :

Sales 80*25 +90*25 = $4,250

Less: COGS = $2,440

Gross Profit = $1,810

3)Moving Average Cost

Average cost is used to account for sales

Transaction

Quantity

Rate

Amount

Jan1 Beginning Inventory

60

15

900

Jan5 Purchases

110

14

1,540

Jan8 Sale

(90)

14.35

(1,291.5)

Jan10 Sales Return

10

14.35

143.5

Jan15 Purchase

35

18

630

Jan16 Purchase Return

(10)

18

(180)

Jan20 Sales

(90)

14.82

(1,333.8)

Jan25 Purchase

10

20

200

Ending Inventory

35

608.2

Cost of Goods Sold = $1,291.5 – 143.5 + 1,333.8 = $2,481.8

Gross Profit :

Sales 80*25 +90*25 = $4,250

Less: COGS = $2,481.8

Gross Profit = $1,768.2